OpenStudy (anonymous):

what happens to the volume of a cone if the radius is not changed but the height is doubled

OpenStudy (anonymous):

V = (1/3) pi r^2 h new volume = (1/3) pi r^2 (2h) = 2 * [(1/3)pi r h] the volume doubles

OpenStudy (anonymous):

i meant 2 * [(1/3)pi r^2 h]

OpenStudy (anonymous):

so it basically is doubling the original answer???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kk thank you so much